OpenStudy (anonymous):

Using the Pythagorean Theorem, if 62 + b2 = 102, then b = ?

hartnn (hartnn):

you mean \(\large 6^2+b^2 =10^2\) right ?

hartnn (hartnn):

subtract 6^2 from both sides, what u get ?

OpenStudy (texaschic101):

Is it 6^2 or 62 ?

OpenStudy (anonymous):

yes that's what I meant @hartnn

OpenStudy (texaschic101):

do you know what 6^2 is ?

OpenStudy (anonymous):

12 @texaschic101

OpenStudy (texaschic101):

no.....6^2 means 6 * 6 and 10^2 = 10 * 10 firgure those out and put them in your equation

OpenStudy (texaschic101):

do you need help with that ?

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

do I do 36*100 ?

OpenStudy (texaschic101):

6^2 + b^2 = 10^2 36 + b^2 = 100 b^2 = 100 - 36 b^2 = 64 b = square root of 64 can you solve for b now ?

OpenStudy (anonymous):

8

OpenStudy (texaschic101):

you got it. Do you have any questions, because I will be happy to answer them

OpenStudy (anonymous):

Describe a way you could draw two lines which are perpendicular to each other without the use of a compass and protractor, using objects you see in the classroom.

OpenStudy (texaschic101):

I am sorry....I do not know this one

OpenStudy (anonymous):

A line that crosses parallel lines but does NOT form a 90 degree angle is called?

OpenStudy (texaschic101):

not sure...I think it is a transversal. You will need a second opinion on this. Geometry is not my best subject.